OpenStudy (anonymous):

I have a distance problem and I know the equation is d=rt. I'm just not sure how to solve the problem: A flight crew flew 420 km in 3 h with a tailwind. Flying against the wind, the flight crew flew 440 km in 4 h. Find the rate of the flight crew in calm air and the rate of the wind

OpenStudy (anonymous):

Well, assume that the tail wind adds some amount (x) to the rate.

OpenStudy (anonymous):

So instead of the normal equation you have d = (r+x)t for flying with a tailwind, and d = (r-x)t for flying against the wind. Then solve for r and x.

OpenStudy (anonymous):

you could try this. the speed with the tail wind is \[\frac{420}{3}=140\] so with the the tailwind they are going 140 kph and against the wind it is \[\frac{440}{4}=110\] kph then as popak said, the first speed going is r + x and the speed returning is r - x so \[r+x=140\] and \[r-x=110\]